# Attention Sinks in Diffusion Language Models

### v1

@@ 5.3 Long Context Modelling
In ARMs, attention sinks have been proven to act as a tool to control over-mixing and avoid representation collapse, especially in long contexts (Barbero et al., 2025; Di Giovanni et al., 2023). However, attention sinks in ARMs are usually present only at the beginning of the sequence and represent a single point of reference for the entire generation. In contrast, DLMs offer a flexible inference and their sinks often shift position during generation. By dynamically directing attention to tokens that are currently most important for the ongoing prediction, DLMs might be able to maintain strong, long-range connections more effectively than ARMs that rely on a single, fixed bottleneck for information. Having the ability to access sinks at the end of the sequence might represent and advantage for long reasoning and planning tasks (Ye et al., 2024, 2025a, 2025b), where the model needs a reference anchor in the future instead of the usual static one at the beginning of the sequence.
@@ 7 Future Work
While our empirical analysis offers a general overview of sink behaviour in DLMs, it also raises several open questions. First, it remains unclear what type of information the model stores in the sinks that correspond to future positions. A promising direction to investigate this would be a mechanistic analysis, for instance using the Logit Lens (Nostalgebraist, 2023).
Second, it is worth exploring whether sinks could be exploited for acceleration or compression, similar to their original use case in (Xiao et al., 2023).
@@ 8 Limitations
While we conducted an extensive study across three DLMs, our analysis is limited to instruct models, as we did not perform experiments on their corresponding base versions. Furthermore, we focused on attention sinks in pre-trained models and did not explore how modifications to the training procedure might influence their behaviour, an aspect that has recently been investigated for ARMs by Miller (2023); OpenAI et al. (2025).
@@ Appendix B Selection of Sink Threshold
In Equation 3 we defined $\epsilon$ to be the threshold for classifying a token as a sink. In Figure˜10 we show how the value of $\epsilon$ affects sink selection. We see that most all the analysed DLMs filter out at least 96% of tokens when usigng $\epsilon=3$.
Figure 10: Percentage of tokens selected when increasing the value of $\epsilon$, for a sequence of 64 tokens. A balanced threshold is found at $\epsilon=3$, which we used in this investigation to define that a token is a sink.

### v2

@@ 5.2 Robustness to Masking Sinks
Model Token Freq.
LLaDA Ċ 0.368
<|mdm_mask|> 0.366
. 0.050
<|start_header_id|> 0.008
(white-space) 0.008
Dream <|mask|> 0.321
Ċ 0.090
. 0.046
, 0.044
<|endoftext|> 0.040
MMaDA Ċ 0.180
<|mdm_mask|> 0.166
<|end_header_id|> 0.160
<|startoftext|> 0.108
istant 0.066
Table 2: Relative Frequency of Top-5 Sink Tokens. Relative frequency of the top-5 most frequent tokens detected as sinks by our metric. Across all models the Ċ (a token representing white-spaces) and [MASK] token are the ones at which the models most often sink their attention. We note that in MMaDA the istant token comes from assistant , which is part of the prompt template. Frequencies computed across 30 generations with generation length of 128, 128 denoising steps and block size of 32.
@@ 5.3 Long Context Modelling
In ARMs, attention sinks have been proven to act as a tool to control over-mixing and avoid representation collapse, especially in long contexts (barbero2025llmsattendtoken; oversquashing). However, attention sinks in ARMs are usually present only at the beginning of the sequence and represent a single point of reference for the entire generation. In contrast, DLMs offer a flexible inference and their sinks often shift position during generation. By dynamically directing attention to tokens that are currently most important for the ongoing prediction, DLMs might be able to maintain strong, long-range connections more effectively than ARMs that rely on a single, fixed bottleneck for information. Having the ability to access sinks at the end of the sequence might represent and advantage for long reasoning and planning tasks (ye2024diffusion; ye2024beyond; ye2025implicit), where the model needs a reference anchor in the future instead of the usual static one at the beginning of the sequence.
@@ 7 Future Work
While our empirical analysis offers a general overview of sink behaviour in DLMs, it also raises several open questions. First, it remains unclear what type of information the model stores in the sinks that correspond to future positions. A promising direction to investigate this would be a mechanistic analysis, for instance using the Logit Lens (logitlens).
Second, it is worth exploring whether sinks could be exploited for acceleration or compression, similar to their original use case in (streamingllm).
@@ 8 Limitations
While we conducted an extensive study across three DLMs, our analysis is limited to instruct models, as we did not perform experiments on their corresponding base versions. Furthermore, we focused on attention sinks in pre-trained models and did not explore how modifications to the training procedure might influence their behaviour, an aspect that has recently been investigated for ARMs by offbyone; gptoss.
@@ 9 Acknowledgements
Figure 9: Distribution of attention scores in Dream-7B and MMaDA-8B
@@ Appendix B Selection of Sink Threshold
In Equation 3 we defined $\epsilon$ to be the threshold for classifying a token as a sink. In Figure˜10 we show how the value of $\epsilon$ affects sink selection. We see that all the analysed DLMs filter out at least 96% of tokens when using $\epsilon=3$.
Figure 10: Percentage of tokens not considered as sinks when increasing the value of $\epsilon$, for a sequence of 64 tokens. A balanced threshold is found at $\epsilon=3$, which we used in this investigation to define that a token is a sink.